# Zero-Order Sharpness-Aware Minimization

### v1

@@ C.1 Additional Details on Synthetic Functions Experiments
$F(\bm{\theta})=\frac{1}{2}\sum_{i=1}^{d}\theta_{i}^{2},$ (108)
$F(\bm{\theta})=\sum_{i=1}^{d}|\theta_{i}|^{3}+\frac{\theta_{i}^{2}}{2},$ (109)
@@
Experimental Setup for Additional Dimensions: To comprehensively evaluate the effectiveness of the zeroth-order (ZO) optimizer across varying problem dimensions, we extended our experimental analysis to include settings with $d=1000$ and $d=5000$, in addition to $d=10000$. This expansion allows us to assess the scalability and robustness of ZOSA under more challenging conditions. For all optimizers, including ZOSA and established baselines, we performed a hyperparameter search for learning rates within the range [1e5,1e2][1e-5,1e-2], ensuring a fair comparison between different dimensionalities and optimization landscapes.
@@
Figure 4: Results on the four test functions with problem dimension $d=1000$ and $K=100$.
@@
Figure 5: Results on the four test functions with problem dimension $d=5000$ and $K=500$.
@@
To further tailor the experimental setup to the unique characteristics of ZOSA, we adjusted the number of perturbation vectors $K$ (equivalent to $m$ in ZOSA) and the smoothing parameter $\mu$ based on the problem dimension. Specifically, we set $K=100$, $K=500$, and $K=1000$ for $d=1000$, $d=5000$, and $d=10000$ respectively, with $\mu=0.005$ across all cases. Additionally, for ZOSA, we performed a specialized hyperparameter search for its parameter $\rho$ (controlling the sharpness-aware perturbation scale) within the range [1e7,1e3][1e-7,1e-3]. This adaptive tuning of $\rho$ allows ZOSA to dynamically adjust its perturbation magnitude, leveraging its variance-aware mechanism to enhance gradient estimation accuracy in high-dimensional spaces.
@@
Fig. 4 and Fig. 5 present the detailed comparison of the performance of all optimizers across these dimensions, highlighting convergence rates and stability. Fig. 7 and Fig. 7 illustrate the cosine similarity between the true gradient and the estimated gradient, a critical metric for evaluating the quality of ZO gradient approximations. The results demonstrate that ZOSA consistently achieves higher cosine similarity values, particularly in high-dimensional and non-convex settings, owing to its innovative use of variance-reduced first moment estimates and refined second moment scaling. This superior alignment with true gradients underscores ZOSA’s advantage in delivering stable and accurate gradient estimates, even as dimensionality increases. Furthermore, the adaptive learning rate adjustment based on $\sigma_{t_{pert}}$ enables ZOSA to outperform baselines by effectively navigating complex optimization landscapes, making it a robust choice for real-world applications.
@@
Figure 6: Cosine similarity between the estimated gradients and the true gradients with problem dimension $d=1000$.
@@
Figure 7: Cosine similarity between the estimated gradients and the true gradients with problem dimension $d=5000$.
@@ C.2 Additional Details on Zero-Order Prompt Fine-tuning
Implementation Details. We employ a fixed, randomly initialized projection matrix $A\in\mathbb{R}^{d\times D}$ to map a vector $v\in\mathbb{R}^{d}$ into the token embedding space $\mathbb{R}^{D}$. Consequently, we focus on optimizing $v\in\mathbb{R}^{d}$ rather than the prompt $p\in\mathbb{R}^{D}$ directly. The pre-trained RoBERTa-large model (Liu et al., 2019) serves as the foundational architecture, with the matrix $A$ sampled from a normal distribution as outlined in (Sun et al., 2023), specifically $\mathcal{N}(0,\sigma_{e}\sqrt{d})$, where $\sigma_{e}$ represents the standard deviation of word embeddings in RoBERTa-large. For ZOSA, along with comparative methods including CMA-ES (Hansen, 2006), MMES (He et al., 2020), BES (Gao & Sener, 2022), INGO (Lyu & Tsang, 2022), and SABO (Ye et al., 2024), the cross-entropy loss on the training data serves as the zero-order optimization objective across six datasets, with optimization of $v$ conducted subject to a budget of 8,000 evaluations. Initial Gaussian distributions are set with mean $\mu_{0}=0$ and covariance $\Sigma_{0}=I$, with a perturbation population size $m$ for ZOSA searched over $\{4,8,16\}$ and $N=100$ for all other optimizers. Hyperparameter tuning via grid search is applied to ZOSA, INGO, BES, and SABO, with the learning rate $\eta$ for ZOSA explored over $\{1e-6,1e-5,1e-4\}$, the sharpness-aware neighborhood size $\rho$ over $\{1e-6,1e-5,1e-4,1e-3,1e-2\}$ for ZOSA and over $\{10,50,100,500\}$ for SABO, the learning rate $\beta$ for INGO, BES, and SABO over $\{0.1,0.5,1,5\}$, and the spacing parameter $c$ for BES over $\{0.1,1,10\}$. Additionally, we evaluate all methods across varying dimensions of $v$, specifically $d\in\{200,500,1000\}$. Each experiment is replicated three times independently, reporting mean objective values with standard deviations.

### v2

@@ C.1 Additional Details on Synthetic Functions Experiments
Experimental Setup for Additional Dimensions: To comprehensively evaluate the effectiveness of the zeroth-order (ZO) optimizer across varying problem dimensions, we extended our experimental analysis to include settings with $d=1,000$ and $d=5,000$, in addition to $d=10,000$. This expansion allows us to assess the scalability and robustness of ZOSA under more challenging conditions. For all optimizers, including ZOSA and established baselines, we performed a hyperparameter search for learning rates within the range $10^{-5},10^{-2}]$, ensuring a fair comparison between different dimensionalities and optimization landscapes.
@@
Figure 4: Results on the four test functions with problem dimension $d=1,000$ and $K=100$.
@@
Figure 5: Results on the four test functions with problem dimension $d=5,000$ and $K=500$.
@@
To further tailor the experimental setup to the unique characteristics of ZOSA, we adjusted the number of perturbation vectors $K$ (equivalent to $m$ in ZOSA) and the smoothing parameter $\mu$ based on the problem dimension. Specifically, we set $K=100$, $K=500$, and $K=1,000$ for $d=1,000$, $d=5,000$, and $d=10,000$ respectively, with $\mu=0.005$ across all cases. Additionally, for ZOSA, we performed a specialized hyperparameter search for its parameter $\rho$ (controlling the sharpness-aware perturbation scale) within the range $10^{-7},10^{-3}]$. This adaptive tuning of $\rho$ allows ZOSA to dynamically adjust its perturbation magnitude, leveraging its variance-aware mechanism to enhance gradient estimation accuracy in high-dimensional spaces.
@@
Fig. 4 and Fig. 5 present the detailed comparison of the performance of all optimizers across these dimensions, highlighting convergence rates and stability. Fig. 7 and Fig. 7 illustrate the cosine similarity between the true gradient and the estimated gradient, a critical metric for evaluating the quality of ZO gradient approximations. The results demonstrate that ZOSA consistently achieves higher cosine similarity values, particularly in high-dimensional and non-convex settings, owing to its innovative use of variance-reduced first moment estimates and refined second moment scaling. This superior alignment with true gradients underscores ZOSA’s advantage in delivering stable and accurate gradient estimates, even as dimensionality increases. Furthermore, the adaptive learning rate adjustment based on $\sigma_{t_{\text{pert}}}$ enables ZOSA to outperform baselines by effectively navigating complex optimization landscapes, making it a robust choice for real-world applications.
@@
Figure 6: Cosine similarity between the estimated gradients and the true gradients with problem dimension $d=1,000$.
@@
Figure 7: Cosine similarity between the estimated gradients and the true gradients with problem dimension $d=5,000$.
We evaluate our proposed ZOSA against state-of-the-art zero-order (ZO) optimization algorithms, MeZO and R-AdaZO, on synthetic benchmark functions, including Quadratic, Cubic, Levy, and Rosenbrock. For MeZO and R-AdaZO, we set the perturbation dimension $K=1000$. Since ZOSA requires approximately twice the query budget per iteration due to its dual-perturbation scheme, we configure its inner-loop iterations $m=500$ to ensure a fair comparison in terms of total queries. Figure 8 plots the suboptimality gap (in log scale) versus the number of queries.
Figure 8: Results on the four test functions with problem dimension $d=1,0000$.
R-AdaZO, as a momentum-based optimizer, exhibits faster convergence across all functions. MeZO converges rapidly on simpler problems but fails to converge on more challenging multimodal landscapes (e.g., Rosenbrock). In contrast, ZOSA achieves superior overall performance, balancing rapid initial progress with robust convergence to near-optimal gaps, particularly on complex functions, highlighting its effectiveness in adaptive perturbation scaling.
To further validate that ZOSA identifies flatter solutions (characterized by smaller Hessian norms, which indicate broader minima and greater robustness to perturbations), we evaluate the Hessian norm at the converged points across the synthetic benchmarks. Table 2 reports these metrics for a fair comparison under equivalent query budgets. For convex functions like Quadratic and Cubic, the distinctions are subtle: all optimizers achieve comparably low Hessian norms, reflecting the inherent smoothness of these landscapes, with minimal gains from ZOSA’s adaptive scaling. In contrast, on non-convex, multimodal problems such as Levy and Rosenbrock, ZOSA outperforms baselines by yielding smaller Hessian norms. This demonstrates ZOSA’s efficacy in navigating rugged terrains to exploit flat regions, thereby enhancing generalization potential in derivative-free settings.
Table 2: Converged Hessian norms and losses on synthetic benchmarks. We report values at convergence under equivalent query budgets(20020000).
Method Hessian norm Loss Steps m/K
Quadratic
MeZO 1 $^{-5} 1010$0 1000
R-AdaZO 1 0.079211 20000 1000
ZOSA 1 0.015748 10000 500
Cubic
MeZO 1.0159 $^{-5} 1010$0 1000
R-AdaZO 1.0253 0.0802 20000 1000
ZOSA 1.0196 0.0251 10000 500
Levy
MeZO 6.3935 5350.0703 10100 1000
R-AdaZO 6.4910 5089.7446 20000 1000
ZOSA 6.3791 5328.3559 10000 500
Rosenbrock
MeZO nan nan 10100 1000
R-AdaZO 302.1303 9698.1153 20000 1000
ZOSA 236.5117 10275.1484 10000 500
Table 3: Performance (%) on SST-2, RTE, MRPC, AG’s News, SNLI and Yelp P. datasets. We report the mean over 3 random seeds. The best result is highlighted in bold.
Methods SST-2 RTE MRPC AG’s News SNLI Yelp P. Average
ZOSA (Ours) 89.11 55.6 81.35 82.66 41.05 91.77 73.59
ZOSA w/o SAM 86.81 50.54 79.64 80.96 38.67 90.69 71.22
ZOSA w/o VAS 85.89 53.43 79.15 75.75 38.82 89.64 70.45
MeZO 86.01 53.43 79.15 78.42 38.83 90.27 71.02
R-AdaZO 88.88 54.51 79.76 82.28 39.06 91.41 72.65
FZOO 88.53 53.43 79.15 81.00 39.65 92.38 72.36
@@ C.2 Additional Details on Zero-Order Prompt Fine-tuning
Implementation Details. We employ a fixed, randomly initialized projection matrix $A\in\mathbb{R}^{d\times D}$ to map a vector $v\in\mathbb{R}^{d}$ into the token embedding space $\mathbb{R}^{D}$. Consequently, we focus on optimizing $v\in\mathbb{R}^{d}$ rather than the prompt $p\in\mathbb{R}^{D}$ directly. The pre-trained RoBERTa-large model (liu2019roberta) serves as the foundational architecture, with the matrix $A$ sampled from a normal distribution as outlined in (sun2023), specifically $\mathcal{N}(0,\sigma_{e}\sqrt{d})$, where $\sigma_{e}$ represents the standard deviation of word embeddings in RoBERTa-large. For ZOSA, along with comparative methods including CMA-ES (hansen2006), MMES (he2020), BES (gao2022), INGO (lyu2022), and SABO (ye2024), the cross-entropy loss on the training data serves as the zero-order optimization objective across six datasets, with optimization of $v$ conducted subject to a budget of 8,000 evaluations. Initial Gaussian distributions are set with mean $\mu_{0}=0$ and covariance $\Sigma_{0}=I$, with a perturbation population size $m$ for ZOSA searched over $\{4,8\}$ and $N=100$ for all other optimizers. Hyperparameter tuning via grid search is applied to ZOSA, INGO, BES, and SABO, with the learning rate $\eta$ for ZOSA explored over $10^{-6},10^{-3}]$, the sharpness-aware neighborhood size $\rho$ over $10^{-6},10^{-1}]$ for ZOSA and over $\{10,50,100,500\}$ for SABO, the learning rate $\beta$ for INGO, BES, and SABO over $\{0.1,0.5,1,5\}$, and the spacing parameter $c$ for BES over $\{0.1,1,10\}$. Additionally, we evaluate all methods across varying dimensions of $v$, specifically $d\in\{200,500,1000\}$. Each experiment is replicated three times independently, reporting mean objective values with standard deviations.
Ablation Experiments. We performed ablation studies on ZOSA and benchmarked it against R-AdaZO, MeZO, and FZOO. The hyperparameters for these baselines were tuned via grid search, with the learning rate $\eta$ in $10^{-1},10^{-5}]$ and other parameters as follows: for MeZO, $\mu\in\{10^{-3},5\times 10^{-3}\}$, $K\in\{4,8,16\}$; for FZOO, $\epsilon\in\{10^{-3},10^{-4}\}$, $N\in\{4,8,16\}$; for R-AdaZO, $\mu=10^{-3}$, $\beta_{1}=0.9$, $\beta_{2}=0.99$, $K\in\{4,8,16\}$. As illustrated in Table 3, ablating either the SAM or VAS (variance-based adaptive scaling) component from ZOSA leads to performance declines across all datasets, with the relative impacts of SAM and VAS varying by dataset. On the five datasets excluding RTE, VAS contributes marginally more to accuracy than SAM; conversely, on the RTE dataset, SAM exhibits a substantially greater contribution than VAS. Nonetheless, removing either VAS or SAM significantly undermines ZOSA’s overall average accuracy. Furthermore, ZOSA surpasses MeZO on all datasets, indicating that batched Rademacher estimation is more appropriate for these six datasets compared to Gaussian perturbations. Relative to R-AdaZO, ZOSA demonstrates modestly superior performance across every dataset.